a cylindrical tank of 2m diameter is laid with its axis horizontal and is filled with water just to the top,the force on one of its end plate in kn is


1
Expert's answer
2021-01-10T18:27:56-0500
"F=mg=\rho g V =0.25\pi d^2 h\rho g \\F =0.25\pi (2)^2 (4)(9.8)(1)=123\ kN"